s - a(subscript 1)*r = a(subscript 2)*r + a(subscript 3)*r ...(Solve for r).
I worked this problem and got r = s/a(subscript 2)*a(subscript 3)*a(subscript 1)*3. This problem is really confusing and the answer is too.

Answer by richwmiller(17219)

I am going to change those subscripts to x,y, and z.
x=a(sub1)
y=a(sub2)
z=a(sub3)
s-x*r=y*r+z*r
let's get all the r's on one side
s=x*r+y*r+z*r
s=r*(x+y+z)
s/(x+y+z)=r
which is what you have.